OpenStudy (lena772):

Would I use point-slope form to solve this? What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3

OpenStudy (anonymous):

first find slope: m = ( y2 - y1) / ( x2 - x1)

OpenStudy (lena772):

8-2/-5--3 6/-2 slope -3

OpenStudy (anonymous):

then use point slope formula: (y - y1) = m (x - x1)

OpenStudy (lena772):

y-2=-3(x--3) y-2=-3x-9

OpenStudy (lena772):

3x+y=-7 so B?

OpenStudy (anonymous):

y + 3x = -7 rewrite it: 3x + y = -7 yes B

OpenStudy (anonymous):

very good ! :)